OpenStudy (anonymous):

Which answer shows this equation in standard form? Medal and fan. 7 – 3(x – y) = –5x + 2

OpenStudy (anonymous):

@hero @mooniepie @ZAO1234

OpenStudy (anonymous):

7 – 3(x – y) = 5x + 2 7 - 3x + 3y = 5x + 2 -8x + 3y = -5 8x - 3y = 5 which is the standard form of the equation

OpenStudy (anonymous):

3y-8x=-5

OpenStudy (anonymous):

wait. The answer A is the same as yours but it has -8 and -5

OpenStudy (anonymous):

ok